OpenStudy (anonymous):

Find the point on the terminal side of θ = pi/4 that has an x coordinate of 2.

OpenStudy (anonymous):

-y = xtan(π/4) = 2(1) = 2 (x,y) = (2,2) --x = (r)cos(θ) 2 = (r)cos(π/4) r = 2√2 y = (2√2)cos(π/4) = 2 The point is (2,2) (r,θ) = (2√2,π/4)

hero (hero):

Or more simply \(y = \large\frac{2}{tan{45^{\circ}}}\)

hero (hero):

Since you already know the x coordinate, you have (2,y)

OpenStudy (anonymous):

so it's 1.23?

hero (hero):

@yahoo! already posted what the point is clear as day.

OpenStudy (anonymous):

OH THANKS VERY MUCH GUYS! i appreciate it! :)